OpenStudy (anonymous):

So confused!!! Can you guys please help!!!! If f(x) varies directly with x2, and f(x) = 75 when x = 5, find the value of f(4). A. 100 B. 50 C. 48 D. 25

OpenStudy (j-cross18):

Ummm ii think it is B but i might be wrong lol

OpenStudy (anonymous):

k

OpenStudy (j-cross18):

cool

OpenStudy (anonymous):

Lol but i need to make sure

OpenStudy (anonymous):

\[f(x)=k\times x^2\] \[f(5)=75=k\times 5^2\] \[75=25k\] find \(k=3\) and so \[f(x)=3\times x^2\]

OpenStudy (anonymous):

then \[f(4)=3\times 4^2=3\times 16\]

OpenStudy (anonymous):

Thanks @satellite73